Let us define a function f : N → Q as follows,

(n-1)/4 n is odd

f(n) = {

(n+1)/2 n is even

Which of the following are true?

f is not one to one.

f is not onto.

f is onto.

f is one to one.


1
Expert's answer
2022-01-05T05:06:35-0500

Let us define a function "f : \N → \mathbb Q" as follows, "f(n) =\begin{cases}
\frac{n-1}4, \text{ if } n \text{ is odd}\\
\frac{n+1}2, \text{ if } n \text{ is even}
\end{cases}."


Since "f(7)=\frac{7-1}4=\frac{6}4=\frac{3}2" and "f(2)=\frac{2+1}2=\frac{3}2," we conclude that "f(7)=f(2)," and hence a function "f : \N → \mathbb Q" is not one-to-one.

Taking into account that "\frac{n-1}4=\frac{1}8" implies "n=\frac{1}2+1=\frac{3}2\notin \N" and "\frac{n+1}2=\frac{1}8" implies "n=\frac{1}4-1=-\frac{3}4\notin \N," we conclude that for "\frac{1}8\in\mathbb Q" the preimage "f^{-1}(\frac{1}8)=\emptyset," and thus a function "f : \N → \mathbb Q" is not onto.


Therefore, the statements ""f" is not one to one" and ""f" is not onto" are true, and the staterments ""f" is onto" and ""f" is one to one" are false.